OpenStudy (anonymous):

Laura is baking a cake. The recipe says that she has to mix 64 grams of chocolate powder to the flour. Laura knows that 1 cup of that particular chocolate powder has a mass of 128 grams. She added two over three of a cup of chocolate powder to the flour. Should Laura add more chocolate powder to make the exact recipe, or did she go over and by what amount?

OpenStudy (anonymous):

So what do you know so far?

OpenStudy (anonymous):

I honestly know nothing.

OpenStudy (anonymous):

@shark35 can you help?

OpenStudy (flashp):

1/6 is the answer

OpenStudy (anonymous):

does she need to add or did she go over by 1/6

OpenStudy (flashp):

She needed to add 64g which is half a cup She added 2/3 She went over by 2/3 - 1/2 = 1/6

OpenStudy (anonymous):

yes @autumnajj

OpenStudy (anonymous):

thanks @flashp

OpenStudy (anonymous):

The length of the base of a triangle is 6.6 cm and its height is 4.9 cm. If the area of the triangle is half of the product of its base length and height, the area of the triangle is __________ cm2. (Round your answer to the appropriate significant digits)

OpenStudy (anonymous):

She needed to add 64g which is half a cup She added 2/3 She went over by 2/3 - 1/2 = 1/6

OpenStudy (anonymous):

for this one The length of the base of a triangle is 6.6 cm and its height is 4.9 cm. If the area of the triangle is half of the product of its base length and height, the area of the triangle is __________ cm2. (Round your answer to the appropriate significant digits)

OpenStudy (anonymous):

1/2 * 6.6 * 4.9

OpenStudy (anonymous):

area for a triangle is 1/2 base * height given base is 6.6 and 4.9 is height

OpenStudy (anonymous):

So how do i put that in numerically? without thestars

OpenStudy (anonymous):

16.7 cm^2

OpenStudy (anonymous):

Ohhhhh so wait, how did you find that?

OpenStudy (anonymous):

Paul plans to put concrete on a rectangular portion of his driveway. The portion is 12 feet long and 6 inches high. The price of concrete is $98 per cubic yard. The total cost of the concrete Paul needs is $108.89. Which of the following is closest to the width of the portion of the driveway on which Paul plans to put concrete? [1 foot = 12inches; 1 yard = 3 feet]

OpenStudy (anonymous):

The price given is per cubic yard, and a cubic yard is defined as the volume of a cube with sides of 1 yard. So, let us convert everything to yards first. 12 feet= 4 yards 6 inches=0.166667 yards Rectangular portion area=4*0.166667*X yards (X is the unknown width) Paul needs a total cost of 108.89 and a cubic yard costs 98. So, how many cubic yards can he cover with that? =108.89/98= 1.1111 cubic yards Hence, 4*0.166667*X=1.1111 X=1.6666 yards which is equal to approximately 5 feet.

OpenStudy (anonymous):

yeah i couldn't figure that out

OpenStudy (anonymous):

Jake cut out 3 small circles and 5 big circles from a piece of cardboard. The radius of each small circle was x inches and the radius of each big circle was y inches. The expression below shows the total number of square inches of cardboard that Jake used for cutting out the circles: 3πx2 + 5πy2 What does the term 5πy2 in the expression represent?

OpenStudy (anonymous):

small circle was x big circle was y so what those 5pi y^2 represents ( the one with the y )

OpenStudy (anonymous):

5 big circles that is the 5 from 5 pi y^2 area of the circle = pi.y^2

OpenStudy (anonymous):

so its the total number of all the circles?

OpenStudy (anonymous):

or total area?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

just to let you know that for the number two, its ^2 not 2

OpenStudy (anonymous):

do you have your answer

OpenStudy (anonymous):

no. is it the total area of all the circles or is it like the total number of circles??

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

im pretty sure its the total area of all the circles

OpenStudy (anonymous):

ooooh okay

OpenStudy (anonymous):

Betty had $40 in her pocket. She bought 12 identical bags of popcorn. Betty was left with $10 after buying the popcorn. What was the price, in dollars, of each bag of popcorn?

OpenStudy (anonymous):

40 -10 = $30. 30/12 = $ 2.50 per bag of popcorn ANSWER

OpenStudy (anonymous):

Evaluate 32 + (5 - 2) • 4 - 6 over 3.

OpenStudy (anonymous):

42 32 + (5 - 2) · 4 - 6 over 3 = 32 + 3x4 - 6/3 =32 + 12 - 2 = 42

OpenStudy (anonymous):

42 isn't one of the options

OpenStudy (anonymous):

19, 16, 46, or 24

OpenStudy (anonymous):

wat are the options

OpenStudy (anonymous):

ok wait

OpenStudy (anonymous):

okay